# revtex4-2Repair the float package

### v1

@@ III.6.2 Sub-threshold Behaviour
Below the code threshold, it is well known that the logical error rate of the code will scale as
@@ IV Discussion and Conclusion
In this paper, we have investigated a scheme for quantum computation in which the binomial code was concatenated with the surface code and realised by MBQC. We subjected this scheme to a realistic model of photon loss and compared several methods of phase measurement and qubit state inference. We also incorporated the soft information available from phase measurements into the decoding process by using a hybrid Dijkstra’s algorithm decoder. Using these techniques, we found phase diagrams of code thresholds as a function of loss, $\gamma$ and measurement error, $q$, for codes of varying orders of discrete rotational symmetry, $N$. Finally, we investigated the performance of these codes in the subthreshold regime.
@@
The other main performance improvement we saw came from using Dijkstra’s algorithm decoding rather than unweighted MWPM decoding. By using the soft information of the continuous variable measurement outcomes, we were able to achieve a threshold against loss of ${}20\%$ , compared to ${}14\%$ using MWPM. This is not only a significant improvement, but is objectively a high threshold against loss.
@@ Appendix C X Basis Pauli Twirl
Suppose that we are using the binning algorithm QSI. There are $2N$ binning regions centered at $pi\cdot l/N$ for $l\in\{0,1,...,2N-1\}.$ Each binning region spans the angles $\pi(l-1/2)/N\rightarrow\pi(l+1/2)/N$ where even $l$ corresponds to $+1$ bins and odd $l$ corresponds to $-1$ bins. The POVM for ‘measuring a ±$\pm 1$ X basis outcome’ is therefore given by the sum over binning regions for even (odd) $l$, respectively. We will now integrate the term with $\phi$ dependence from Eq. LABEL:unsimplified, $e^{i(k-k^{\prime})N\phi},$ over the binning regions

### v2

@@ III.6.2 Sub-threshold Behaviour
Below the code threshold, it is well known [42] that the logical error rate of the code will scale as
@@ IV Discussion and Conclusion
In this paper, we have investigated a scheme for quantum computation in which the binomial code was concatenated with the surface code and realised by MBQC. We subjected this scheme to a realistic model of photon loss and compared several methods of phase measurement and qubit state inference. We also incorporated the soft information available from phase measurements into the decoding process by using a hybrid modified MWPM algorithm decoder. Using these techniques, we found phase diagrams of code thresholds as a function of loss, $\gamma$ and measurement error, $q$, for codes of varying orders of discrete rotational symmetry, $N$. Finally, we investigated the performance of these codes in the subthreshold regime.
@@
The other main performance improvement we saw came from using the modified algorithm decoding rather than unweighted MWPM decoding. By using the soft information of the continuous variable measurement outcomes, we were able to achieve a threshold against loss of ${}20\%$ , compared to ${}14\%$ using MWPM. This is not only a significant improvement, but is objectively a high threshold against loss.
@@ Appendix C X Basis Pauli Twirl
Suppose that we are using the binning algorithm QSI. There are $2N$ binning regions centered at $\pi\cdot l/N$ for $l\in\{0,1,...,2N-1\}.$ Each binning region spans the angles $\pi(l-1/2)/N\rightarrow\pi(l+1/2)/N$ where even $l$ corresponds to $+1$ bins and odd $l$ corresponds to $-1$ bins. The POVM for ‘measuring a ±$\pm 1$ X basis outcome’ is therefore given by the sum over binning regions for even (odd) $l$, respectively. We will now integrate the term with $\phi$ dependence from Eq. 62, $e^{i(k-k^{\prime})N\phi},$ over the binning regions